# AceFF: A State-of-the-Art Machine Learning Potential for Small Molecules

### v1

@@ Abstract
We introduce AceFF, a pre-trained machine learning interatomic potential (MLIP) optimized for small molecule drug discovery. While MLIPs have emerged as efficient alternatives to Density Functional Theory (DFT), generalizability across diverse chemical spaces remains difficult. AceFF addresses this via a refined TensorNet2 architecture trained on a comprehensive dataset of drug-like compounds. This approach yields a force field that balances high-throughput inference speed with DFT-level accuracy. AceFF fully supports the essential medicinal chemistry elements (H, B, C, N, O, F, Si, P, S, Cl, Br, I) and is explicitly trained to handle charged states. Validation against rigorous benchmarks, including complex torsional energy scans, molecular dynamics trajectories, batched minimizations, and forces and anergy accuracy demonstrates that AceFF establishes a new state-of-the-art for organic molecules. The AceFF-2 model weights and inference code are available at https://huggingface.co/Acellera/AceFF-2.0.
@@ 1 Introduction
Atomistic simulations are fundamental tools for understanding materials and molecular systems at the atomic level. The accuracy and utility of these simulations critically depend on the choice of force field. Classical molecular mechanics (MM) force fields such as GAFFwang2004development, wang2006automatic, CGenFFvanommeslaeghe2010charmm, vanommeslaeghe2012automation, OpenFFqiu2021development, boothroyd2023development, and AMBERmaier2015ff14sb offer high computational efficiency, but often fall short in predictive accuracy, particularly for diverse, drug-like molecules containing rare functional groups or where quantum effects and polarization play significant roles. First-principles quantum mechanical (QM) methods, such as density functional theory (DFT), provide substantially higher accuracy but are prohibitively expensive for routine biomolecular simulations.
@@
Machine learning interatomic potentials (MLIPs) offer a promising alternative, achieving near-QM accuracy at a fraction of the computational cost: an order of magnitude slower than MM methods, while being many orders of magnitude faster than QM approaches Behler2007NNP, Bartok2010GAP, Smith2017ANI1, Unke2021MLFFReview. The current state-of-the-art MLIP architectures in terms of accuracy include equivariant message-passing graph neural network models such as MACEbatatia_mace_2023, Nequipbatzner_e3-equivariant_2022, AIMNet2 anstine2024aimnet2, and TensorNetTensorNet, to name a fewduval2023hitchhiker. One of the most used models to date is ANI-2xANI2x, which employs a less accurate invariant architecture but is one of the most computationally efficient MLIPs available. A newer and more accurate group of pre-trained models are the MACE-OFF23 models with large, medium, and small variants available. These are significantly more accurate than ANI-2x, with the downside of being substantially slower kovacs2023mace. Other pre-trained models include: AIMNet2anstine2024aimnet2 which supports charged molecules, AceFF-1.0aceff_huggingface based on the TensorNet architecturesimeon_broadening_2025 supporting charge and spins, the Egret models, which are based on the MACE architecturemann_egret-1_2025; the recent OrbMolorbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024; and UMA-OMol25 modelslevine2025openmolecules2025omol25. OrbMol and UMA are both trained on the recent OMol25 dataset levine2025openmolecules2025omol25 and cover most of the periodic table, while being accurate in standard benchmarks, they are slow relative to more specialized models.
@@
It is currently not feasible to perform production MD using a full MLIP description of the entire molecular system. Instead, it is necessary to use hybrid MLIP/MM schemes, where only the ligand is modeled with the MLIP, and the remainder of the system uses a classical MM force field. It has been found that even with a simple mechanical embedding approach, MLIP/MM can provide increased accuracy for RBFE calculationssabanes2024enhancing, sabanes_zariquiey_quantumbind-rbfe_2025. In our previous work, we were limited by the fact that ANI-2x supports only eight elements and only neutral molecules. This led us to create our own QM dataset and develop our first MLIP, AceFF-1.0aceff_huggingface, which we applied in a large-scale RBFE benchmark, finding improved accuracy compared to traditional force fields and previous ANI-2x resultssabanes_zariquiey_quantumbind-rbfe_2025. AceFF-1.1 included a larger dataset and was released on HuggingFaceaceff1.1_huggingface. Both of these models are freely available under the Apache 2.0 license for academic and commercial use. In this paper, we report the latest update to our model family, AceFF-2, which includes both more training data and an improved model architecture, TensorNet2. Additionally, we describe the various benchmarks we use to quickly assess model accuracy, making comparisons in accuracy with the existing MLIPs ANI-2x ANI2x, AIMNet2 anstine2024aimnet2, OrbMol orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale; semi-empirical methods g-XTB froitzheim_g-xtb_2025, GFN2-XTB bannwarth_gfn2-xtbaccurate_2019; and the MM forcefield GAFF2 wang2004development, wang2006automatic. In summary, AceFF-2 offers accuracy higher than even slower models, reaching state-of-the-art on standard benchmarks, at speed. Because of this, we believe that it has immediate applications in drug discovery.
@@ 2.1 TensorNet2 architecture
AceFF-1.x models were based on TensorNetTensorNet models with charge and spin encodingssimeon_broadening_2025. With AceFF-2, we introduce a modified version of TensorNet with improved charge handling. The approach we take closely follows that of AIMNet2anstine2024aimnet2: we include additional scalar features to mimic partial charges, perform neutral charge equilibration on them, and use them in a long-range Coulomb energy term. We name this version TensorNet2. An overview of the model architecture is shown in figure 1.
@@
Figure 1: The AceFF-2 architecture TensorNet2 builds upon TensorNetTensorNet with the new components colored in blue. The main change is that after the tensor embedding, and each tensor interaction block, a set of partial charges $q$ and weights $w$ are computed and undergo a neutral charge equilibration (NQE) procedure as done in AIMNet2anstine2024aimnet2. The partial charges are then combined with edge features during the tensor interaction. The predicted charges are used to compute a Coulomb energy $E_{\text{Coulomb}}$ alongside the standard short range node-wise MLIP energy term $E_{i}$. The dashed arrow represents where iteration occurs for more than one interaction layer. The pictured diagram corresponds to TensorNet2 1-layer. The grey dashed circles around the example molecule illustrate the short cutoffs used in the tensor embedding and interaction. The blue dashed circle represents the long range of the total charge (a global feature feature), NQE block, and the Coulomb interaction.
@@
From an input of atomic numbers and coordinates $Z_{i}$$R_{i}$, TensorNet learns a set of 3x3 matrices (rank-2 tensors) for each atom $X^{(i)}$ which are equivariant to rotation and translation. At each message-passing step, the $X^{(i)}$ are updated based on the neighboring atoms, making use of the fact that $X^{(i)}$ can be decomposed into scalar, vector, and symmetric traceless components, $I$, AA, $S$, respectively. After several message passing updates, the squared Frobenius norms of the representations, which are invariant under O(3), are further processed by the neural network to predict energies, obtaining atomic forces via automatic differentiation. Moreover, a distinctive feature of the TensorNet message-passing architecture, specifically when interaction layers are active, is its natural transition from O(3) to SO(3) equivariance. This reduction in symmetry follows directly from how tensor–tensor interaction products are computed in the interaction layer, which mix even and odd parities and therefore break reflection equivariance while preserving full rotational equivariance. This behavior is especially valuable in drug discovery, where chirality is central: enantiomers are related by reflections but often differ dramatically in binding affinity and biological activityh2011significance.
@@
In the previous version of TensorNetsimeon_broadening_2025, total molecular charge $Q$ was included in a simple manner with no extra learned parameters. While this method effectively resolved the degeneracy problem for molecules sharing the same coordinates and atomic numbers but differing in $Q$, we observed limitations during extrapolation. Specifically, when tested on a set of larger charged molecules, the performance was significantly surpassed by the AIMNet2 architecture anstine2024aimnet2, which explicitly models and learns partial charges and Coulomb interactions.
@@
which is the same as in anstine2024aimnet2. The $q_{i}$ from each node are then used in the next message passing update by extending the scalar edge features such that Equation 11 in TensorNet becomes:
@@
This procedure constitutes a type of self-consistent method and is fundamentally different in style from alternative charge inclusion techniques such as the latent Ewald summation method cheng_latent_2025, which calculates an Ewald summation or a Coulomb-like interaction using node features derived from just the final node features. Another is the 4G-HDNNP approach ko_fourth-generation_2021, which employs neural networks to predict electronegativity and hardness. This prediction is followed by a full charge equilibration scheme, which involves solving a set of linear equations to find the set of partial charges that minimizes the electrostatic energy subject to the total charge constraint, which scales poorly with system size. The TensorNet2 neutral charge equilibration approach adds only a small overhead to training and inference speed. For example, a 1-layer TensorNet model with an embedding dimension of 128 and 32 radial basis functions has 535,681 parameters, while a 1-layer TensorNet2 model with the same embedding dimension, radial basis functions, and additionally a charge dimension of 16 has 685,413 parameters. The inference speeds of each model for a 1500 atom system (500 water molecules) are 35 steps/second and 32 steps/second, respectively. A parameter increase of 28% and a speed decrease of under 10%.
As part of the TensorNet2 development, we optimized the tensor embedding and tensor interaction steps of both TensorNet and TensorNet2, resulting in a speed increase of approximately 30% and a reduction in memory usage of approximately 30%. The key change was to maintain $I$, AA, and $S$ as scalar, vector, and tensor-shaped arrays rather than the $3\times 3$ tensors used in the original code. We also re-implemented the custom CUDA neighbor list in Triton to ease development and deployment, also achieving a minor speed increase. For example, for 1,000 atoms, the Triton neighbor list code has a latency of 0.0136 ms, while the CUDA code has a latency of 0.0194 ms. It should be noted that a full TensorNet evaluation has a latency of at least 10 ms for 1,000 atoms; therefore, the neighbor list speeds and differences between them are negligible (see Methods).
@@ 2.2 AceFF-2 trained model
We trained a 2-layer version of TensorNet2 on an internal QM dataset. We used a charge embedding dimension of 16, a hidden dimension of 128, and 32 radial basis functions. The model was trained using the torchmd-train scripts that can be found in TorchMD-Net pelaez2024torchmd. We refer to this model as AceFF-2. The model was trained on energy and force labels, without using any partial charge or dipole moment data. Optionally, one of the TensorNet2 charge channels could be trained on a QM level specific type of partial charge, e.g. MBIS, to facilitate an approach such as semelak_advancing_2025. Indeed, we expect the different charge channels could be trained simultaneously on multiple QM-derived partial charges—for example, the Omol25 dataset contains Mulliken and Lowdin charges levine2025openmolecules2025omol25. For ablation purposes, we also trained a 1-layer version with all other hyperparameters unchanged.
@@ 2.3 Accuracy on torsion scans
The first benchmark test we performed was Sellers et al torsion scan sellers2017comparison. This benchmark data set contains torsion scans of 62 molecules done with different levels of QM. We take the most accurate method from the set, CCSD(T)/CBS, and use this as the reference to compare against. For all molecules in the test, we performed relaxed torsion scans (as done inSmith2017ANI1, smith_transforming_2018). For comparison we did the same procedure with our AceFF-1.0, ANI-2xANI2x from torchanigao2020torchani, AIMNet2anstine2024aimnet2, and the recent OrbMol-conservativeorbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024. We also did the scan with the semi-empirical method GFN2-XTBbannwarth2019gfn2 and the molecular mechanics force-field GAFFwang2004development, wang2006automatic. We excluded g-XTB because it does not have analytical gradients available yet. The MAE of each torsion scan relative to the reference scan is plotted in 2a as box-plots with all data points shown.
Figure 2: Torsion scans. a. Sellers et al. torsion scan benchmarksellers2017comparison. The orange lines are the median value. The methods are ordered from left to right by median MAE compared to the coupled cluster reference data. The y-axis has been intentionally capped at 2.0 kcal/mol to focus on the differences between the MLIP models. An image of one of the molecules with the torsion angle indicated is pictured. Values marked by * were taken from a previous publication sabanes_zariquiey_quantumbind-rbfe_2025. All other data points were recalculated. b. Behara et al torsion scan benchmark behara2024. The orange lines are the median value. The methods are ordered from left to right by median MAE compared to the coupled cluster reference data. The scatter points are color-coded by the molecular charge. The y-axis has been intentionally capped at 4.0kcal/mol to focus on the differences between the MLIPs. Data points of methods marked with * were taken from the original databehara2024, all other data points were recalculated. On both plots, a guide line has been drawn at 1 kcal/mol.
@@
We also performed a second torsion scan benchmark from Behara et al. behara2024. This torsion scan contains charged molecules, specifically [0,-1,+1]. The reference torsion drives were done with MP2/heavy-aug-cc-pVTZ, and the reference single point energies were labeled with CCSD(T)/CBS. The benchmark differs from the Sellers et al torsion scan in that we do not perform any geometry optimizations. This benchmark tests only the energy error of the MLIP and not the force calculations. We did the single point calculations for OrbMol, AceFF-2, AceFF-1.0, AIMNet2, GFN2-XTB, and GAFF. We took the single point results for MACE-OFF23 and the DFT method wB97M-V/dzvp frombehara2024 for extra comparison.
[…]

### v2

@@ Abstract
We introduce AceFF, a pre-trained machine learning interatomic potential (MLIP) optimized for small molecule drug discovery. While MLIPs have emerged as efficient alternatives to Density Functional Theory (DFT), generalizability across diverse chemical spaces remains difficult. AceFF addresses this via a refined TensorNet2 architecture trained on a comprehensive dataset of drug-like compounds. This approach yields a force field that balances high-throughput inference speed with DFT-level accuracy. AceFF fully supports the essential medicinal chemistry elements (H, B, C, N, O, F, Si, P, S, Cl, Br, I) and is explicitly trained to handle charged states. Validation against rigorous benchmarks, including complex torsional energy scans, molecular dynamics trajectories, batched minimizations, and tests of force and energy accuracy, demonstrates that AceFF is state-of-the-art for organic molecules in the accuracy and speed regime important for drug discovery. The AceFF-2 model weights and inference code are available at https://huggingface.co/Acellera/AceFF-2.0.
@@ 1 Introduction
Atomistic simulations are fundamental tools for understanding materials and molecular systems at the atomic level. The accuracy and utility of these simulations critically depend on the choice of force field. Classical molecular mechanics (MM) force fields such as GAFF24, 23, CGenFF21, 22, OpenFF15, 4, and AMBER13 offer high computational efficiency, but often fall short in predictive accuracy, particularly for diverse, drug-like molecules containing rare functional groups or where quantum effects and polarization play significant roles. First-principles quantum mechanical (QM) methods, such as density functional theory (DFT), provide substantially higher accuracy but are prohibitively expensive for routine biomolecular simulations.
@@
Machine learning interatomic potentials (MLIPs) offer a promising alternative, achieving near-QM accuracy at a fraction of the computational cost: an order of magnitude slower than MM methods, while being many orders of magnitude faster than QM approaches Behler2007NNP, Bartok2010GAP, Smith2017ANI1, Unke2021MLFFReview. The current state-of-the-art MLIP architectures in terms of accuracy include equivariant message-passing graph neural network models such as MACEbatatia_mace_2023, Nequipbatzner_e3-equivariant_2022, AIMNet2 2, and TensorNet20, to name a few7. One of the most used models to date is ANI-2x5, which employs a less accurate invariant architecture but is one of the most computationally efficient MLIPs available. A newer and more accurate group of pre-trained models are the MACE-OFF23 models with large, medium, and small variants available. These are significantly more accurate than ANI-2x, with the downside of being substantially slower 12. Other pre-trained models include: AIMNet22 which supports charged molecules, AceFF-1.01 based on the TensorNet architecturesimeon_broadening_2025 supporting charge and spins, the Egret models, which are based on the MACE architecturemann_egret-1_2025; the recent OrbMolorbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024; and UMA-OMol25 modelslevine2025openmolecules2025omol25. OrbMol and UMA are both trained on the recent OMol25 dataset levine2025openmolecules2025omol25 and cover most of the periodic table, while being accurate in standard benchmarks, they are slow relative to more specialized models. A recent fast model is SO3LR kabylda_molecular_2025 which incorporates semi-local interactions from a equivariant neural network with universal pairwise force fields designed for short-range repulsion, long-range electrostatics, and dispersion interactions.
@@
It is currently not feasible to perform production MD using a full MLIP description of the entire molecular system. Instead, it is necessary to use hybrid MLIP/MM schemes, where only the ligand is modeled with the MLIP, and the remainder of the system uses a classical MM force field. It has been found that even with a simple mechanical embedding approach, MLIP/MM can provide increased accuracy for RBFE calculations18, sabanes_zariquiey_quantumbind-rbfe_2025. In our previous work, we were limited by the fact that ANI-2x supports only eight elements and only neutral molecules. This led us to create our own QM dataset and develop our first MLIP, AceFF-1.01, which we applied in a large-scale RBFE benchmark, finding improved accuracy compared to traditional force fields and previous ANI-2x resultssabanes_zariquiey_quantumbind-rbfe_2025. AceFF-1.1 included a larger dataset and was released on HuggingFaceaceff1.1_huggingface. Both of these models are freely available under the Apache 2.0 license for academic and commercial use. In this paper, we report the latest update to our model family, AceFF-2, which includes both more training data and an improved model architecture, TensorNet2, and a faster implementation with optimized Nvidia warp kernels. Additionally, we describe the various benchmarks we use to quickly assess model accuracy, making comparisons in accuracy with the existing MLIPs ANI-2x 5, AIMNet2 2, OrbMol orbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, UMA levine2025openmolecules2025omol25; semi-empirical methods g-XTB froitzheim_g-xtb_2025, GFN2-XTB bannwarth_gfn2-xtbaccurate_2019; and the MM forcefield GAFF2 24, 23. In summary, AceFF-2 is one of the fastest models with high accuracy, because of this, we believe that it has immediate applications in drug discovery.
@@ 2.1 TensorNet2 architecture
AceFF-1.x models were based on TensorNet20 models with charge and spin encodingssimeon_broadening_2025. With AceFF-2, we introduce a modified version of TensorNet with improved charge handling. The approach we take closely follows that of AIMNet22: we include additional scalar features to mimic partial charges, perform neutral charge equilibration on them, and use them in a long-range Coulomb energy term. We name this version TensorNet2. An overview of the model architecture is shown in figure 1.
@@
Figure 1: The AceFF-2 architecture TensorNet2 builds upon TensorNet20 with the new components colored in blue. The main change is that after the tensor embedding, and each tensor interaction block, a set of partial charges $q$ and weights $w$ are computed and undergo a neutral charge equilibration (NQE) procedure as done in AIMNet22. The partial charges are then combined with edge features during the tensor interaction. The predicted charges are used to compute a Coulomb energy $E_{\text{Coulomb}}$ alongside the standard short range node-wise MLIP energy term $E_{i}$. The dashed arrow represents where iteration occurs for more than one interaction layer. The pictured diagram corresponds to TensorNet2 1-layer. The grey dashed circles around the example molecule illustrate the short cutoffs used in the tensor embedding and interaction. The blue dashed circle represents the long range of the total charge (a global feature feature), NQE block, and the Coulomb interaction.
@@
From an input of atomic numbers and coordinates $Z_{i}$$R_{i}$, TensorNet learns a set of 3x3 matrices (rank-2 tensors) for each atom $X^{(i)}$ which are equivariant to rotation and translation. At each message-passing step, the $X^{(i)}$ are updated based on the neighboring atoms, making use of the fact that $X^{(i)}$ can be decomposed into scalar, vector, and symmetric traceless components, $I$, AA, $S$, respectively. After several message passing updates, the squared Frobenius norms of the representations, which are invariant under O(3), are further processed by the neural network to predict energies, obtaining atomic forces via automatic differentiation. Moreover, a distinctive feature of the TensorNet message-passing architecture, specifically when interaction layers are active, is its natural transition from O(3) to SO(3) equivariance. This reduction in symmetry follows directly from how tensor–tensor interaction products are computed in the interaction layer, which mix even and odd parities and therefore break reflection equivariance while preserving full rotational equivariance. This behavior is especially valuable in drug discovery, where chirality is central: enantiomers are related by reflections but often differ dramatically in binding affinity and biological activity11.
@@
In the previous version of TensorNetsimeon_broadening_2025, total molecular charge $Q$ was included in a simple manner with no extra learned parameters. While this method effectively resolved the degeneracy problem for molecules sharing the same coordinates and atomic numbers but differing in $Q$, we observed limitations during extrapolation. Specifically, when tested on a set of larger charged molecules, the performance was significantly surpassed by the AIMNet2 architecture 2, which explicitly models and learns partial charges and Coulomb interactions.
@@
which is the same as in 2. The $q_{i}$ from each node are then used in the next message passing update by extending the scalar edge features such that Equation 11 in 20 becomes:
@@
This procedure constitutes a type of self-consistent method and is fundamentally different in style from alternative charge inclusion techniques such as the latent Ewald summation method cheng_latent_2025, which calculates an Ewald summation or a Coulomb-like interaction using node features derived from just the final node features. Another is the 4G-HDNNP approach ko_fourth-generation_2021, which employs neural networks to predict electronegativity and hardness. This prediction is followed by a full charge equilibration scheme, which involves solving a set of linear equations to find the set of partial charges that minimizes the electrostatic energy subject to the total charge constraint, which scales poorly with system size. The TensorNet2 neutral charge equilibration approach adds only a small overhead to training and inference speed. For example, a 1-layer TensorNet model with an embedding dimension of 128 and 32 radial basis functions has 535,681 parameters, while a 1-layer TensorNet2 model with the same embedding dimension, radial basis functions, and additionally a charge dimension of 16 has 685,413 parameters. The inference speeds of each model for a 1500 atom system (500 water molecules) are 100 steps/second and 75 steps/second, respectively. A parameter increase of 28% and a speed decrease of 25%.
2.2 Software Optimizations
With this model update, we have introduced software optimizations to the TensorNet models within the TorchMD-Net14 package. These optimizations utilize custom NVIDIA warp kernels, implemented from the MatGL packageKo2025, matgl_repo, for the primary bottleneck operations: the decomposition and composition between the $X$ and $I,A,S$ tensors, as well as the message-passing routines. This implementation results in a 3x speed-up for both inference and training. Figure 2 illustrates the molecular dynamics inference speed (time per step) as a function of system size, comparing the legacy code to the optimized version.
Figure 2: Molecular dynamics speed (time per step) for increasing number of atoms. The blue line is the old pure PyTorch version of TensorNet and the orange line is the new version using the Warp kernels available in TorchMD-Net version 3. These timings are both for the same TensorNet 1 layer model
The updated code also achieves a 3x reduction in GPU memory usage. These improvements stem from a more efficient representation of the $I,A,S$ tensors. The original implementation stored these as three separate $3\times 3$ tensors to facilitate PyTorch vectorization. Mathematically, however, $I,A,$ and $S$ represent a scalar, a vector, and a symmetric traceless tensor, which can be represented by 1, 3, and 5 values, respectively. The new optimized kernels maintain this compact format throughout, eliminating the overhead of expanding to $3\times 3$ tensors populated with zero entries. Furthermore, the message passing kernels avoid the creation of large edge sized intermediate tensors are required for Autograd.
@@ 2.3 AceFF-2 trained model
We trained a 2-layer version of TensorNet2 on an internal QM dataset. We used a charge embedding dimension of 16, a hidden dimension of 128, and 32 radial basis functions. The model was trained using the torchmd-train scripts that can be found in TorchMD-Net 14. We refer to this model as AceFF-2. The model was trained on energy and force labels, without using any partial charge or dipole moment data. Optionally, one of the TensorNet2 charge channels could be trained on a QM level specific type of partial charge, e.g. MBIS, to facilitate an approach such as semelak_advancing_2025. Indeed, we expect the different charge channels could be trained simultaneously on multiple QM-derived partial charges—for example, the Omol25 dataset contains Mulliken and Lowdin charges levine2025openmolecules2025omol25. For ablation purposes, we also trained a 1-layer version with all other hyperparameters unchanged.
@@ 2.4 Accuracy on torsion scans
The first benchmark test we performed was Sellers et al torsion scan 19. This benchmark data set contains torsion scans of 62 molecules done with different levels of QM. We take the most accurate method from the set, CCSD(T)/CBS, and use this as the reference to compare against. For all molecules in the test, we performed relaxed torsion scans (as done insmith_approaching_2019, smith_transforming_2018). For comparison we did the same procedure with our AceFF-1.0, ANI-2x5 from torchani10, AIMNet22, and the recent OrbMol-conservativeorbmol_huggingface, rhodes2025orbv3atomisticsimulationscale, neumann_orb_2024. We also did the scan with the semi-empirical method GFN2-XTB3 and the molecular mechanics force-field GAFF24, 23. We excluded g-XTB because it does not have analytical gradients available yet. The MAE of each torsion scan relative to the reference scan is plotted in 3a as box-plots with all data points shown.
Figure 3: Torsion scans. a. Sellers et al. torsion scan benchmark19. The orange lines are the median value. The methods are ordered from left to right by median MAE compared to the coupled cluster reference data. The y-axis has been intentionally capped at 2.0 kcal/mol to focus on the differences between the MLIP models. An image of one of the molecules with the torsion angle indicated is pictured. Values marked by * were taken from a previous publication sabanes_zariquiey_quantumbind-rbfe_2025. All other data points were recalculated. b. Behara et al torsion scan benchmark behara2024. The orange lines are the median value. The methods are ordered from left to right by median MAE compared to the coupled cluster reference data. The scatter points are color-coded by the molecular charge. The y-axis has been intentionally capped at 4.0kcal/mol to focus on the differences between the MLIPs. Data points of methods marked with * were taken from the original databehara2024, all other data points were recalculated. On both plots, a guide line has been drawn at 1 kcal/mol.
@@
We also performed a second torsion scan benchmark from Behara et al. behara2024. This torsion scan contains charged molecules, specifically [0,-1,+1]. The reference torsion drives were done with MP2/heavy-aug-cc-pVTZ, and the reference single point energies were labeled with CCSD(T)/CBS. The benchmark differs from the Sellers et al torsion scan in that we do not perform any geometry optimizations. This benchmark tests only the energy error of the MLIP and not the force calculations. We did the single point calculations for OrbMol, AceFF-2, AceFF-1.0, AIMNet2, GFN2-XTB, and GAFF. We took the single point results for MACE-OFF23 frombehara2024 for extra comparison.
The MAE of each method is plotted in figure 3b. The trend between models is approximately the same as the Sellers et al torsion scan. As expected, the biggest outliers in AceFF-1.0 are the charged molecules, and similarly for MACE-OFF23 which was designed for only neutral molecules. While UMA, OrbMol, and AceFF2 demonstrate a more balanced performance across both neutral and charged molecules, charged species still represent the most significant deviations for all three. Notably, AIMNet2 is the only model where the largest outlier is a neutral molecule.